OpenStudy (anonymous):

HELP... what is the slope of this line -16x+12y=0

OpenStudy (abbles):

First you need to put it into Y intercept form.. are you familiar with this? y = mx + b

OpenStudy (smored):

Ax+By=C slope=-A/B

OpenStudy (abbles):

Start by isolating the y. -16x + 12y = 0 12y = 16x y = 16/12x y = 4/3x

OpenStudy (abbles):

And since x is the slope, the slope of your line is 4/3.

OpenStudy (anonymous):

4/3

OpenStudy (smored):

\[Ax+By=C\]so A=-16, B=12, and C=0\[-16x+12y=0\]\[12y=16x+0\]\[y=\frac{4}{3}x\]now it is in the slope-intercept form, y=mx+b, and the slope is m. slo slope=4/3